Project

General

Profile

16-May-2016

Attending: Bruno, Chris, Vlad, Joseph, Alex R., Satish, Adam

Status and plans

We're at the very tail end of SA. Upcoming there is a clean up of FD data, CAF respin, maybe another one, due to veto not being stored. CAF respin for the weekend. Focus mostly turning into additional samples.

Computing issues

Kanika opened a ticket with issues about dCache. Concat went from hours to days in ETA, because they’re transfer-time limited.

Samples

FD RHC

Adam has non swap and flux FCLs ok, of which non swap 149 left in the draining, and flux swap 743 missing. Definitions created for the reconstruction. He’ll start this afternoon. RHC appears twice on the name, so Adam will send an email with Joseph in CC. Adam will start reco this afternoon.

ND with high intensity

Adam is trying to figure out how to overlay the rock twice.

FD data closure

The beam database has finally been fixed, so Cristiana will run on the draining definitions.

MRE systematics

Alex has completed 3 out of 5 samples. Caleb x and y systematics are still missing.

Nue-veto cosmics

99% done, so we called it finished.

CAF respins

100% done by Kanika

Concats

100% done, waiting for data. They still show up as being owned by anorman, because the fix is not in the trunk. Kanika has also observed errors in about 1/50 of the jobs, not coming back and failing to copy back.

AOB

Dataset snapshotting is going to happen today, Alex is making a script. We would also like to make a tool that checks the number of events in a file, essentially to check that the concats are produced fine.